# Dynamical Evolutions in Globular Clusters and Dwarf Galaxies: Conduction Fluid Simulations

### v1

@@ I Introduction
The remainder of the paper is organized as follows: We sketch our simulation method in Section II and validate it in Section III. In Section IV, we simulate a system with stars and sub-solar-mass or sup-solar-mass DM particles, comparing our results to N-body simulations in the literature. We also present simulations for various parameters of stellar-DM systems. In Section V, we apply our simulation results to the observed GCs and dwarfs in the Local Volume and provide a preliminary answer to what likely happens to DM in GCs, assuming that significant DM may have been present at the formation of such systems (which is not fully resolved by early universe simulations). We conclude in Section VI. Appendix A provides further details of our algorithm for heat conduction and dynamical heating or cooling, showing that it always satisfies the Courant-Friedrichs-Lewy (CFL) condition for numerical stability. Appendix B gives the translation of fiducial units from ours to those adopted in the literature. Appendix C compares the results from the fluid model and those from the FP model.
@@ II.2 The Coulomb logarithms
in the simulations below. Given this hierarchy, the difference in $\gamma$ should have a minor effect on the simulation results. Therefore, we will fix $\gamma=0.11$ in the computation of the Coulomb logarithms (Eq. (10)) below. We also note that Ref. 1992MNRAS.257..513H ; Graham:2023unf ; Graham:2024hah impose different choices for the argument of the Coulomb logarithms, which again will not significantly change the results because of the slow-varying property of the logarithm function. Note that Eq. (11) implies
@@ II.3 Two-fluid equations
    $\frac{\partial M_{l}}{\partial r}={}$ $4\pi r^{2}\rho_{l},$ (14)
@@
    $\frac{\partial}{\partial r}\left(\rho_{l}\sigma_{l}^{2}\right)={}$ $-\frac{G(M_{l}+M_{h})\rho_{l}}{r^{2}},$ (16)
@@
    Energy Conservation: For the heavier component:
@@
    where $D/Dt$ is the Lagrangian time derivative. Note that the first and second terms on the right-hand sides of Eqs. (17) and (18) are the gravothermal conduction and dynamical heating/cooling exchange terms, respectively.
@@ II.6 Numerical Implementation
To implement the two-fluid equations outlined earlier, we first convert the physical quantities $x=(r,\rho,M,\sigma,t,L)$ into dimensionless quantities $\hat{x}\equiv x/x_{0}$ by introducing a set of fiducial quantities $x_{0}$. The fiducial quantities are constructed based on the scale density and radius of the initial profile of one fluid; here, we choose those of the heavier fluid:
@@
where $c_{1}=(3\pi)^{-1/2}=0.326$, and $c_{2}=2^{1/2}\alpha b/9=0.086$. Note that in deriving Eq. (42) from Eqs. (17) and (18), we have assumed that the densities, $\rho_{l}$ and $\rho_{h}$, remain unchanged during the heat conduction and dynamical heating/cooling step. They will be changed and the velocity updated during the relaxation step, thereby implementing an “operator splitting” technique (see more details below). Given the mass hierarchy Eq. (11), Eqs. (42) and (43) can be simplifies into
@@
where $F$ represents the discretized expression of the right-hand side of Eq. (49) and the supscript on the variable represents the number of time steps.
@@ IV.1 Comparing the two-fluid simulation and the N-body simulation
$t_{\text{fric}}=0.416\frac{M_{\text{tot}}^{1/2}r_{50\%\star}^{3/2}}{G^{1/2}m_{\star}\ln\Lambda_{\star}}=1.08\sqrt{\frac{M_{\text{tot}}}{M_{\text{tot}\star}}}\frac{\ln\Lambda_{\star\chi}}{\ln\Lambda_{\star}}t_{0}.$ (55)
@@ IV.2 Two-fluid simulations in the limit $m_{h}\gg m_{l}$
Ultra-heavy DM candidates, such as PBHs, are well-motivated and can be effectively probed via astrophysical observations. As a specific example, constraints from microlensing, gravitational waves, and galactic dynamics have ruled out PBHs as the 100% of DM for $m\gtrsim 10^{-10}\,\text{M}_{\odot}$. Nevertheless, there are still mass windows where PBHs can contribute a non-negligible fraction of the DM. For instance, around $m\simeq 10^{-6}\,\text{M}_{\odot}$, PBHs can make up $\sim 10\%$ of DM, and around $m\simeq 20\,\text{M}_{\odot}$, they can similarly account for $\sim 10\%$. See the review by Carr:2021bzv .
Here, we are interested in considering a progenitor galaxy made with stars and PBHs in one of the mass windows mentioned above. In either case, we can treat the mass of the heavier component to be much greater than the lighter one, e.g., Eq. (11). Given the hierarchy, we consider nine cases with different scale density ratios ($\xi=0.1,1,10$) and different scale radius ratios ($\zeta=0.5,1,2$) of the two fluids.
@@ V Comparing the simulation results to the observed data for GCs and dwarf galaxies
Given the simulation results presented in Section III.1 and Section IV.2, it is reasonable to adopt the fiducial time scale of the heavier component, $t_{0}^{h}$, defined in Eq. (37), as an approximate indicator of both the mass‑segregation and core‑collapse times for a broad range of $M_{l}/M_{h}$ ratios. [^2] We add a supscript $h$ to emphasize that $t_{0}$ is evaluated for the heavier component. We therefore compare $t_{0}^{h}$ with the ages of GCs and dwarf galaxies in the Local Volume, which are usually estimated to be between $10$ and $13\,\mathrm{Gyr}$. In the following estimate, we set the mass of an individual star to $m_{\star}=1\,\text{M}_{\odot}$ for all GCs or dwarfs. When the DM particle mass satisfies $m_{\chi}\ll m_{\star}$, the typical dynamical timescales of the stellar–DM systems are
@@
For the total stellar mass $M_{\text{tot}\star}$, we will assume $M_{\text{tot}\star}/M_{\text{tot}}\approx 1$ for the GCs. For the dwarf galaxies, we use mass_stellar provided by LVDB, which computes the stellar mass from the V-band absolute magnitude, assuming a mass-to-light ratio of 2. To estimate the total mass $M_{\text{tot}}$, we use the dynamical mass $M_{\text{dyn}}$, defined as the total mass within the half-mass radius. Therefore $M_{\text{tot}}=2M_{\text{dyn}}$. It has been shown that, for a wide range of density profiles, $M_{\text{dyn}}$ is well-approximated by Wolf:2009tu
@@
The resulting scatter plot of $M_{\text{dyn}}$ vs. $t_{0}^{h}$ is shown as the left panel of Fig. 7. We use red dots to represent GCs that appear to have core collapsed, [^5] Observationally, collapsed status is often identified by checking if the density profile of the system cannot be well fitted by a Plummer profile, or any other cored density profile 2005ApJS..161..304M . Here, we adopted the status provided by Harris GC catalog 1996AJ….112.1487H ; Harris:2010ut . blue stars to represent GCs that have no core collapsed, and black squares to represent the dwarfs. The horizontal band indicates the typical age of the systems, $t_{\text{age}}=10-13$ Gyr. As the figure shows, most of the GCs are around the collapsed phase $t_{0}^{h}\simeq t_{\text{age}}$; this includes all the GCs that have been observed to have core collapsed. On the other hand, the majority of dwarf galaxies have $t_{0}^{h}$ much greater than $t_{\text{age}}$, and thus they show almost no core collapse (manifested as dense stellar cores) or mass segregation. Exceptions include those dwarfs with the smallest stellar masses and stellar half-mass radii.
@@ Acknowledgements
We thank Anirudh Chiti, Liang Dai, Alex Drlica-Wagner, Alexander Ji, Andrey Kravtsov, Burcin Mutlu-Pakdil, Ethan Nadler, and Kaixiang Wang for useful discussions. The simulations were partly performed at the University of Chicago’s Research Computing Center. YZ acknowledges the Aspen Center for Physics, which is supported by NSF Grant PHY-1607611, and the Kavli Institute for Theoretical Physics, which is supported by NSF grant NSF PHY-2309135, for their hospitality during the completion of this study. YZ is supported by the GRF Grant 11302824 from the Research Grants Council of Hong Kong SAR. SLS acknowledges support from NSF PHY 2006066 and PHY 2308242 to the University of Illinois at Urbana-Champaign.
@@ A.1 A semi-implicit algorithm to solve the conduction/interaction equation
where the subscript $j$ is the zone index, and the supscript $n$ is the time step index (should not be confused with power). To get the value of $\left(\sqrt{\hat{u}_{h}}\right)^{n+1}_{j}$, we apply an Taylor expansion around $\sqrt{\hat{u}}$ at $n$-th step,
@@ A.2 Checking the CFL condition
for arbitrary choice of $\Delta\ln\hat{r}$ and $\Delta\hat{t}$. Therefore, the algorithm is unconditionally stable.

### v2

@@ I Introduction
The remainder of the paper is organized as follows. We sketch our simulation method in Section II and validate it in Section III. In Section IV, we simulate a system with stars and sub-solar-mass or sup-solar-mass DM particles, comparing our results to N-body simulations in the literature. We also present simulations for various parameters of stellar-DM systems. In Section V, we apply our simulation results to the observed GCs and dwarfs in the Local Volume and provide a preliminary answer to what likely happens to DM in GCs, assuming that significant DM may have been present at the formation of such systems (which is not fully resolved by early universe simulations). We conclude in Section VI. Appendix A provides further details of our algorithm for heat conduction and dynamical heating or cooling, showing that it always satisfies the Courant-Friedrichs-Lewy (CFL) condition for numerical stability. Appendix B gives the translation of fiducial units from ours to those adopted in the literature. Appendix C compares the results from the fluid model and those from the FP model.
@@ II.2 The Coulomb logarithms
in the simulations below. Given this hierarchy, the difference in $\gamma$ should have a minor effect on the simulation results. Therefore, we will fix $\gamma=0.11$ in the computation of the Coulomb logarithms (Eq. (10)) below. We also note that Refs. 1992MNRAS.257..513H ; Graham:2023unf ; Graham:2024hah impose different choices for the argument of the Coulomb logarithms, which again will not significantly change the results because of the slow-varying property of the logarithm function. Note that Eq. (11) implies
@@ II.3 Two-fluid equations
    $\frac{\partial M_{l}}{\partial r}={}$ $4\pi r^{2}\rho_{l};$ (14)
@@
    $\frac{\partial}{\partial r}\left(\rho_{l}\sigma_{l}^{2}\right)={}$ $-\frac{G(M_{l}+M_{h})\rho_{l}}{r^{2}};$ (16)
@@
    Energy Conservation: For the heavier component,
@@
    where $D/Dt$ is the Lagrangian time derivative 1992MNRAS.257..513H . Note that the first and second terms on the right-hand sides of Eqs. (17) and (18) are the gravothermal conduction and dynamical heating/cooling exchange terms, respectively;
@@ II.6 Numerical Implementation
To implement the two-fluid equations outlined earlier, we first convert the physical quantities $x=(r,\rho,M,\sigma,u\equiv 3\sigma^{2}/2,t,L)$ into dimensionless quantities $\hat{x}\equiv x/x_{0}$ by introducing a set of fiducial quantities $x_{0}$. The fiducial quantities are constructed based on the scale density and radius of the initial profile of one fluid; here, we choose those of the heavier fluid:
@@
where $c_{1}=(3\pi)^{-1/2}=0.326$, and $c_{2}=2^{1/2}\alpha b/9=0.086$. Note that in deriving Eq. (42) from Eqs. (17) and (18), we have assumed that the densities, $\rho_{l}$ and $\rho_{h}$, remain unchanged during the heat conduction and dynamical heating/cooling step. They will be changed and the velocity updated during the relaxation step, thereby implementing an “operator splitting” technique (see more details below). Given the mass hierarchy Eq. (11), Eqs. (42) and (43) can be simplified into
@@
where $F$ represents the discretized expression of the right-hand side of Eq. (49) and the superscript on the variable represents the number of time steps.
@@ IV.1 Comparing the two-fluid simulation and the N-body simulation
$t_{\text{fric}}=3.01\sqrt{\frac{M_{\text{tot}}}{M_{\text{tot}\star}}}t_{rh\star}=1.08\sqrt{\frac{M_{\text{tot}}}{M_{\text{tot}\star}}}\frac{\ln\Lambda_{\star\chi}}{\ln\Lambda_{\star}}t_{0}.$ (55)
@@ IV.2 Two-fluid simulations in the limit $m_{h}\gg m_{l}$
Ultra-heavy DM candidates, such as PBHs, are well-motivated and can be effectively probed via astrophysical observations. As a specific example, constraints from microlensing, gravitational waves, and galactic dynamics have ruled out PBHs as 100% of DM for $m\gtrsim 10^{-10}\,\text{M}_{\odot}$. Nevertheless, there are still mass windows where PBHs can contribute a non-negligible fraction of the DM. For instance, around $m\simeq 10^{-6}\,\text{M}_{\odot}$, PBHs can make up $\sim 10\%$ of DM, and around $m\simeq 20\,\text{M}_{\odot}$, they can similarly account for $\sim 10\%$. See the review by Carr:2021bzv .
Here, we are interested in considering a progenitor galaxy made with stars and PBHs in one of the mass windows mentioned above. In either case, we can treat the mass of the heavier component to be much greater than that of the lighter one, e.g., Eq. (11). Given the hierarchy, we consider nine cases with different scale density ratios ($\xi=0.1,1,10$) and different scale radius ratios ($\zeta=0.5,1,2$) of the two fluids.
@@ V Comparing the simulation results to the observed data for GCs and dwarf galaxies
Given the simulation results presented in Section III.1 and Section IV.2, it is reasonable to adopt the fiducial timescale of the heavier component, $t_{0}^{h}$, defined in Eq. (37), as an approximate indicator of both the mass‑segregation and core‑collapse times for a broad range of $M_{l}/M_{h}$ ratios. [^2] We add a superscript $h$ to emphasize that $t_{0}$ is evaluated for the heavier component. We therefore compare $t_{0}^{h}$ with the ages of GCs and dwarf galaxies in the Local Volume, which are usually estimated to be between $10\,\mathrm{Gyr}$ and $13\,\mathrm{Gyr}$. In the following estimate, we set the mass of an individual star to $m_{\star}=1\,\text{M}_{\odot}$ for all GCs or dwarfs. When the DM particle mass satisfies $m_{\chi}\ll m_{\star}$, the typical dynamical timescales of the stellar–DM systems are
@@
For the total stellar mass $M_{\text{tot}\star}$, we will assume $M_{\text{tot}\star}/M_{\text{tot}}\approx 1$ for the GCs. For the dwarf galaxies, we use mass_stellar provided by LVDB, which computes the stellar mass from the V-band absolute magnitude, assuming a mass-to-light ratio of 2. To estimate the total mass $M_{\text{tot}}$, we use the dynamical mass $M_{\text{dyn}}$, defined as the total mass within the half-light radius, and estimate $M_{\text{tot}}\approx 2M_{\text{dyn}}$. It has been shown that, for a wide range of density profiles, $M_{\text{dyn}}$ is well-approximated by Wolf:2009tu
@@
The resulting scatter plot of $M_{\text{dyn}}$ vs. $t_{0}^{h}$ is shown as the left panel of Fig. 7. We use red dots to represent GCs that appear to have core collapsed, [^5] Observationally, collapsed status is often identified by checking if the density profile of the system cannot be well fitted by a Plummer profile, or any other cored density profile 2005ApJS..161..304M . Here, we adopted the status provided by Harris GC catalog 1996AJ….112.1487H ; Harris:2010ut . blue stars to represent GCs that have not core collapsed, and black squares to represent the dwarfs. The horizontal band indicates the typical age of the systems, $t_{\text{age}}=10-13$ Gyr. As the figure shows, most of the GCs are around the collapsed phase $t_{0}^{h}\simeq t_{\text{age}}$; this includes all the GCs that have been observed to have core collapsed. On the other hand, the majority of dwarf galaxies have $t_{0}^{h}$ much greater than $t_{\text{age}}$, and thus they show almost no core collapse (manifested as dense stellar cores) or mass segregation. Exceptions include those dwarfs with the smallest stellar masses and stellar half-mass radii.
@@ Acknowledgements
We thank Anirudh Chiti, Liang Dai, Ariane Dekker, Alex Drlica-Wagner, Moritz Fischer, Alexander Ji, Andrey Kravtsov, Burcin Mutlu-Pakdil, Ethan Nadler, Kaixiang Wang, and Hai-Bo Yu for useful discussions. The simulations were partly performed at the University of Chicago’s Research Computing Center. YZ acknowledges the Aspen Center for Physics, which is supported by NSF Grant No. PHY-1607611, and the Kavli Institute for Theoretical Physics, which is supported by NSF Grant No. PHY-2309135, for their hospitality during the completion of this study. YZ is supported by the GRF Grants No. 11302824 and No. 11310925 from the Research Grants Council of Hong Kong SAR and the Grant No. 9610645 from the City University of Hong Kong. SLS acknowledges support from NSF Grants No. PHY 2006066 and No. PHY 2308242 to the University of Illinois at Urbana-Champaign.
@@ A.1 A semi-implicit algorithm to solve the conduction/interaction equation
where the subscript $j$ is the zone index, and the superscript $n$ is the time step index (should not be confused with power). To get the value of $\left(\sqrt{\hat{u}_{h}}\right)^{n+1}_{j}$, we apply an Taylor expansion around $\sqrt{\hat{u}}$ at $n$-th step,
@@ A.2 Checking the CFL condition
for an arbitrary choice of $\Delta\ln\hat{r}$ and $\Delta\hat{t}$. Therefore, the algorithm is unconditionally stable.